Home‎ > ‎Reports‎ > ‎Crew101‎ > ‎

day08


Crew 101 Log Book for March 19, 2011

Photos of the day





Commander's Check-In Report

Tom Caillouet Reporting

Crew Health: Good.

EVAs and Times: 1350 - 1517

Number of Reports: 5 (Commander, Engineering, Journalist, Chef, EVA)

Field Results: Preliminary recon of radio ridge trail, gather three samples (2 for bio & geo, 1 just for geo).

Plans for Tomorrow: EVAs originally planned for today but canceled due to high wind: re-test radio signal north of the Hab along Cow Dung Road, raise radio repeater antenna. May also include another scouting mission.


Commander's Report

Tom Caillouet Reporting

'Mars' hasn't been the friendliest place while we've been here. Today,  we were planning on having several EVAs to test and improve radio coverage. Unfortunately, high winds canceled these plans for us. Instead, we decided to run another biological sample return EVA, again attempting to go to Cow Dung Reservoir.

I was hoping that having been there yesterday, it would be fairly trivial to find the location. It was decidedly not trivial, though partially this was due to an exploratory urge. We decided to take a branch off the main trail and explore the route to Radio Ridge.

After this detour we collected a sample at a place that looked a bit like the reservoir from the road. We'll give it another shot in another days, but at least we had the total success of pancakes this morning.


Engineering Report

Cody Hall Reporting

Generator/Electricity:

• Kitty (Diesel Generator) (Used? Yes/No): Yes

• Honey (Gas Generator) (Used? Yes/No): No

• Battery Charge Level:

o V everstart

o V armor plate

o V green

Notes/Comments (include how many generator runs, approximately when,

and any times you needed to turn the generator on/off):

1 generator run


Fuel Status:

• Diesel (Full/0.75/0.5/0.25/Empty): 0.75

• Propane Tank (Full/0.75/0.5/0.25/Empty): 0.25

• Gasoline Tank (Full/0.75/0.5/0.25/Empty): temperature too low

to check after use, should be close to 0.75 still

Notes/Comments:


Oil:

• Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

• Honey Oil Quantity (# of Quarts in storage): Leave it to DG

• ATV Oil Quantity (# of Quarts in storage): 6 quarts on the

engineering work bench

• Car Oil Quantity (# of Quarts in storage): Leave it to DG

Notes/Comments:


Water Status:

• Outside Potable Water Tank Level (inches from bottom): 15.5

• Trailer Potable Water Tank Level (inches from bottom): 6

• Hab Potable Water Tank Level (inches from bottom): 12

• Potable Water Meter Reading: 436825

Notes/Comments :


GreenHab:

• Condition of plants in Tank 1: unchanged

• Condition of plants in Tank 2: more duckweed

• Condition of plants in Tank 3: water level dropped 0.25 inches

• Greenhab notes/ comments:


Observatory:

• Telescope Used? (Yes/no) No

Observatory Notes:


Transportation:


Hab Car (NOT RENTAL)

• Used (yes/no): No

• Oil Checked (yes/no): No

• Tire Status: Not checked

• Odometer at end of day: 123688.4

• Notes/Comments on Hab Car:


ATV 1 (Opportunity):

• Used: yes

• Oil Checked: no

• Fuel Consumed: replaced

• Tire Status: ok

• Comments:


ATV 2 (Spirit):

• Used: yes

• Oil Checked: no

• Fuel Consumed: replaced

• Tire Status: ok

• Comments:


ATV 3 (Viking I):

• Used: no

• Oil Checked: no

• Fuel Consumed: none

• Tire Status: ok

• Comments


ATV 4 (Viking II):

• Used: no

• Oil Checked: no

• Fuel Consumed: none

• Tire Status: ok

• Comments:

ATV Comments:


Heating and Ventilation:

• Thermostat setting upstairs (Farenheit): 68

• Actual temperature upstairs (Farenheit): 68

• Thermostat setting downstairs (Farenheit): 62

• Actual temperature downstairs (Farenheit): 61

Notes/Comments:


Computers/Networking Infrastructure: working

Notes/comments on internet/computers:


General Engineering

Notes/Comments:


Science Report

Christina Graves Reporting


On 3-19-11, an ATV EVA (EVA #9) was performed by Commander Caillouet and Christina graves.  The purpose of the EVA was to collect soil samples and geological data from multiple sites on the San Rafael Swell (the site of the MDRS) for analysis.


The San Rafael Swell is characterized by both anticline and sedimentary rocks.  Geological data obtained includes rock morphology, type, and location, as well as Geiger counter readings.


Data obtained from the EVA is found below: 

Site 1

Local time: 14:42

GPS Location Data (UTM):  12S 0517686    4254622    

Accuracy (m): 17

Alt (m): 1399

# of rocks larger than 6 in: 0

depth of collection: 6 in. 

Site description: dried up river beds

Weight of sample: 51.40g


Site 2

Local time: 14:56

GPS Location Data (UTM): Unavailable

# of rocks larger than 6in.:  3

depth of collection: 0.5 in.

Site description: top of hill sample site - pebble collection

Weight of sample: 222.66g


Site 3

Local time: 15:08 

GPS Location Data (UTM): 12S 0510979     4251416

Accuracy (m): 7

Altitude (m): 1378

# of rocks larger than 6 in.: 0

depth of collection: 1 in. 

Site description: large "field" of pebbles - appear green from above. Pebble layer atop red clay.

Weight of sample: 139.69g


Commander Caillouet and crew member Graves arrived at Site 1 at 14:42.  A 6' x 6' area was marked and photographed at 12S 0517686 4254622 UTM (+/- 17m) at an altitude of 1399.  A 51.40g soil sample was collected at a depth of 6.0 in.  The area was purposefully chosen, and was a dried up river bed.  The sample consisted of uniform light brown, fine grit  sand. The site serves as a potential model of a "primordial pool," thought to be necessary for life in prebiotic situations.  There were no larger rocks in the area (larger rocks were classified as those greater than 6 in. in diameter).


The team arrived at Site 2 at 14:56 local time.  Commander Caillouet and Crew member Graves climbed to the top of a hill to obtain a sample collection.  Due to accessibility of the hill, a 6' x 6' plot was not marked as at the other 2 sites. 


The team members attempted to take a GPS location reading (by taking a photograph of the GPS monitor), but the camera was focused improperly, and therefore coordinates were not obtained.  The 222.66g sample consists of varying types of rocks and pebbles, collected at a depth of 0.5 in. There were three large rocks (larger than 6 in. in diameter) in the area. It should be noted that the site was photographed.


The team arrived at Site 3 at 15:08 local time.  A 6' x 6' area was marked and photographed, with GPS coordinates at 12S 0510979 4251416 UTM (+/- 7m) at an altitude of 1378.  The area was appeared as a large "field" of pebbles - appearing green from a distance.  Beneath the layer of small pebbles (no rocks exceeding 0.5 in.) was red clay.  A 139.69g sample was obtained at the site at a depth of 1in. consisting of both pebbles and red clay.


Biological soil analysis (EZ N.A. Soil Analysis from Omega and PCR) will be performed at a later time to test for the presence of bacteria.


EVA 9 Report

Christina Graves Reporting

EVA Time: 1350-1514

EVA Crewmembers: Tommy KJ4KAC, Christina

Base Communications: Jon KF7GLD

EVA Procedure: The crew set out to find Cow Dung Reservoir for soil samples for geology and biology experiments, but were again (as on 3-18-11) unable to locate the site. Soil samples were taken from alternative locations, instead. The sites were chosen based on characteristics of water flow, including at the base of a dried up river bed.

General site locations were north and north east of the Hab. The team was not to deviate from established ATV paths, and ATV helmets were exchanged with bubble helmets prior to sample collection and on-foot exploration of sites.


EVA Results: Soil samples were returned from three sites located at 12S 0517686, 4254622 UTM (Site 1) and 12S 0510979, 4251416 UTM (Site 3). GPS data was not recorded at Site 2 (an attempt was made, but the focus of the camera taking the picture of the GPS was too blurry to get a clear reading). Accuracy of the first GPS location (Site 1) was 17m, with an altitude of 1399m; accuracy of the third GPS location (Site 3) was 7 with an altitude of 1378.

Lessons Learned: Crew members should take water along on longer EVA's.


Chef's Report

Cody Hall Reporting

1a) Today was a cooking day.

1b) Was today a special day celebrated at the main meal with special

food or activity?

No

2) List the foods served at today's main meal (usually the evening

meal), giving full names of each.

Macaroni and Cheese

Chocolate Chip and Walnut Cookies

Bacon

3) List any main meal foods not finished by the crew (leftovers stored

for later use or discarded)

4) Recipes for anything more complicated than following package

directions exactly, or rehydrating.

Macaroni and Cheese

Boil noodles, strain water, pour into oven pan

Add 1 whole egg (powdered), chopped up velveeta (1/3 block, or as

desired), butter (as desired)

Fill half of pan with milk

Bake in oven at 350 degrees for 45 minutes

Remove from oven, stir, put back in oven for a few minutes

Serve

Chocolate Chip and Walnut Cookies (Chrissy made these!)

2 ¼ cups all purpose flour

¾ brown sugar

¾ sugar

1 cup Crisco

¾ cup chips

¾ cup walnuts

1 teaspoon vanilla extract

1 teaspoon salt

1 teaspoon baking soda

2 large eggs

Mix all together and bake at 375 for 16 minutes

5) Lessons learned (if any)

Don’t add too much milk, and add salt and pepper to noodles

6) Comments/questions for Kim and Jean, the food study investigators

We need more butter, I couldn’t add very much because we were almost out.



Journalist Report

Christine Redmond Reporting

Today was another windy day on Mars! This morning I had a little piece of heaven, a warm shower! Oh the smell of Oawis soap! Tommy made pancakes to make the morning an unforgettable one for crew 101!

On a more serious note, despite the windy conditions, Tommy and Christina had another successful EVA during which they collected more biology samples. Jon had a busy afternoon working on the radios. We have been taking full advantage of today's “cook day”! This afternoon I brought home back to the Hab with some toll house cookies! Cody is currently preparing our dinner, mac and cheese

We are all looking forward to another day on Mars!